OpenStudy (_malice_):

A triangular piece of rubber is stretched equally on all three sides (without distorting or changing its shape) such that the new sides are 2 times the length of the original sides. How does the area of the triangle change? A. The area increases to 2 times the original value. B. The area increases to 4 times the original value. C. The area increases to 6 times the original value. D. The area increases to 8 times the original value.

OpenStudy (mathmate):

Hints: Here the scale factor is 2 because new sides are 2 times the lengths of the original sides. Length of each side is measured in metres, so linear dimensions (lengths) vary with scale factor. Area is measured in metre\(^2\), so areas vary with scale factor\(^2\). Volume is measured in metre\(^3\), so volumes vary with scale factor\(^3\).

OpenStudy (_malice_):

Oooh Okay thank you. I gtg now but thank you a ton

OpenStudy (kainui):

area of triangle with base length b and height a: \[A=\frac{ab}{2}\] now let's call the new area the stretched area A prime with a little mark on it like this A'. That means we doubled the lengths: \[A' = \frac{(2*a)(2*b)}{2} = 4\frac{ab}{2}=4A\] Maybe complicated, but what I showed is that the stretched area A' is equal to four times the old area A. Think about it a bit, it's pretty powerful stuff.